Which equation represents a line parallel to the line whose equation is 2y − 5x = 10?
Answer by stanbon(75887)

Which equation represents a line parallel to the line
whose equation is 2y − 5x = 10?
-------------------------------------
slope-intercept form: y = (5/2)x + 5
slope = 5/2
-----------------
An equation with the same slope will be parallel
to the given line is if it does not have a
y-intercept of 5.
=============================
Example:
y = (5/2)x + 7 is parallel to the given line.
==========================
Cheers,
Stan H.